7.6(3)(1 point) Consider the Initial Value Problem -L* 4)*, x0=[!] (a) Find the eigenvalues and eigenvectors for the coefficient ma
Answer #1

- -5 -47x 210) c) | 4 -5 for eigen volver A- Arco 1-5rd Ylo T Y 5.11=0 (5-4)2 + 1600 25+ A 2 + 10x41650 ² +10 4+415o AE -10 =No eigen veckor in ) - :) - To so solution it x17) - eest{ [ Jae- *** + [1] 2 elit? given xlo) :) - 13). ©{ [4] + [:16} MB] -X1H) - / 3-31 )e-st fil6944-i Siny!) + ( Gs 41+ isin 4) 1 1 (Csut - isinut) til Gyt tisina) / +182) [ astf(i+1 (Gs 4t + Singh

X1H) - / 3-31 )e-st fil6944-i Siny!) + ( Gs 41+ isin 4) 1 1 (Csut - isinut) til Gyt tisina) / +182) [ astf(i+1 (Gs 4t + Singh) 7 litid caht- & singh) | lais est s ces 4rt sinut 7 [ Gyt-singh X(t) = [ 3e St ( Ces 4f + sinyt). 3 e-St (es 48-sinut) Answer